1 molal aq. Solution of trichloroacetic and (ccl3cooH) is heated to its boiling point.solution has boiling point of 100.18*C.determine the vant hoff factor for trichloroacetic acid.(kb for water=0.512K kg mol

Grade:12

1 Answers

Naga Himanshu
7 years ago
We know that, Tb = Kb * molality * vant hoff factor0.18 = 0.512 * 1 * ii = 0.35Vant hoff factor = 0.35We can calculate degree of dissociation `x` by i = 1 + (n-1)x